OpenStudy (anonymous):

Which strategies could be used to solve 32 × 49? Choose all answers that are correct. A. (30 + 49) + (2 + 49) = B. (30 × 49) + (2 × 49) = C. (50 × 32) − (50 × 1) = D. (50 × 32) − (1 × 32) = there is more than 1 answer to this

OpenStudy (anonymous):

@jhawes

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (anonymous):

needs to equal 1568 A= 130 (30 + 49 = 79 and 2 + 49 = 51) NO A B = 1568 (30 x 49 = 1470 and 2 x 49 = 98) YES B C= 1550 (50 x 32= 1600 and 50 x 1 = 50 so 1600-50= 1550) so NOT C D= 1568 (50 × 32= 1600) − (1 × 32=32) =1568) YES D So answers are B & D